A metal wire is 3.0 m long at 0 degree Celcius. Its length changes by 0.3 mm when its at 200 degree celcius. Its thermal coefficient is
Answer #1

let, thermal coefficient be c
given,
delta L = 0.3 mm
= (0.3*10^-3) m

use,
(delta L)/L = c*(delta T)
(0.3*10^-3)/3.0 = c*(200 - 0)
10^-4 = c*200
c = (5*10^-7) (degree celcius)^-1

Answer: (5*10^-7) (degree celcius)^-1